Joe drew 32 pictures in his sketch book. He drew four times as many pictures as Kevin. How many pictures did Kevin draw? A train leaves Rock City at an average speed of 50 miles per hour and heads for Gnome City. Another train leaves Gnome City at an average speed of 40 miles per hour and heads for Rock City. If the route is 360 miles long, how many hours will it take for the 2 trains to meet? Alicia was paid $125 for babysitting five days after school for the Smith family. Each day Mrs. Smith paid her $3 more than the day before. How much money did she earn on the first day?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum ー E ゴWrap Text aCopy B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 conditional Format . Cell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells AG1 AB AC AD AE AF AG AH Al AK AL AM 2 ANS 1 JOE DREW 32 PICTURES = 4 TIMES X NO OF PICTURES DREW BY KEVIN SO KEVIN DREW32/4 8 PICTURES 4 ANS 2 6 5 AS THE TRAINS ARE IN OPPOSITE DIRECTION, WE HAVE TO MAKE ADDITION OF SPEEDS OF TWO TRAINS SO TOTAL SPEED = 50 + 40 = 90 THAT IS BOTH TRAINS WILL COVER 90 MILES IN HOUR ROUTE IS 360 MILES LONG SO NO OF HOURS360/90 4 HOURS TO MEET 9 10 11 ANS3 12 13 14 15 16 17 18 19 SUPPOSE ALICIA PAID X ON FIRST DAY SO NEXT DAY SHE WILL PAID X +3, THEN X +3 +3 X +6 AND SO ON 125PAID FOR 5 DAYS 125- X(X +3)(X +6) (X 9) +(X+12) 125-5(X) +30 95 = 5X S 19 PAYMENT ON FIRST DAY HMassignment DEAR Sheet2 COV TIME SERIES Corr REGRESSION CAMERAEXP RETURN MATRD-INTERVAL, NORMAL HYPOTHESIS GPI MATHIS, Sh 07:33